OpenStudy (anonymous):

There is a small error. what's the error ? integral [tan(x) = sin(x)/cos(x)]

OpenStudy (anonymous):

@tkhunny can u solve this

OpenStudy (tkhunny):

My first impression is that the notation is meaningless. Can you formulate a better question?

OpenStudy (anonymous):

\[\int\limits (\tan(x)) dx = \int\limits (\frac{ \sin(x) }{ \cos(x) })\]

OpenStudy (anonymous):

@tkhunny there' error converting this integral...it is small ut what is it

OpenStudy (tkhunny):

Tough call. The little 'dx' is really superfluous if the context is properly understood. More than the 'dx' is the poor LaTeX coding, creating the scrawny parentheses on the right.

OpenStudy (anonymous):

can iu answer it or not

OpenStudy (anonymous):

what level math is it?

OpenStudy (anonymous):

calculus bc

OpenStudy (anonymous):

let me see if I understand the question. you are saying from Tan(x) to Sin(x)/cos(x) there is a error

OpenStudy (anonymous):

ya but there is integral on both side

OpenStudy (anonymous):

are you missing dx in right side on purpose?

OpenStudy (anonymous):

an error exists in the logic that says the \[\int\limits \tan x dx = \int\limits \frac{ \sin x }{ \cos x }\]. explain where and why the error occurs and provide graphical support for your explanation

OpenStudy (anonymous):

here's the question again

OpenStudy (anonymous):

oh there is dx on both side...MY BAD

OpenStudy (anonymous):

well, there is constant in both function , they could be different

OpenStudy (anonymous):

no theres not

OpenStudy (anonymous):

no when you integrate both sides , you get - ln|cos x| + C=? -ln{cos}+D they are equal only when C are D equal

OpenStudy (anonymous):

read the question i posted...the question means to ask something else

OpenStudy (anonymous):

thts not the answer i already verified that with my teacher...answer is something else

OpenStudy (anonymous):

ok , I don't know

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@jamesJ